# Estimating the Shannon Entropy Using the Pitman–Yor Process

### v1

@@ 1 Introduction
In this study, we consider the problem of estimating $H$ in the setting where both $K$ and $\bm{p}$ are unknown, as in [1, 13]. Let $T$ denote the number of observed species in the random sample and let $n_{1},\ldots,n_{T}$ denote their observed frequencies. When the sample size NN is sufficiently large relative to $K$, it is reasonable to assume that $T=K$. In this case, the simplest and most widely used approach is to estimate $p_{1},\ldots,p_{T}$ by maximum likelihood and substitute the resulting estimates into (1.1). Specifically, the maximum likelihood estimator of each probability mass is given by
@@
In the literature, alternative estimators have been proposed, including the Miller–Madow estimator [18] and the Chao–Shen estimator [7]. However, it has been indicated that these methods tend to perform poorly when the sample size NN is not sufficiently large relative to the true number of species $K$ [20, 24]. To address this issue, Hausser and Strimmer [12] proposed a shrinkage estimator for the individual probability masses toward $1/K$ under the assumption that $K$ is known, and demonstrated that the resulting estimator exhibits favorable properties even for small sample sizes. In contrast, we consider the more challenging setting in which the sample size NN is smaller than the true number of species $K$, and $K$ is unknown. Under this regime, we propose a new estimator of $H$. The proposed approach is based on the Pitman–Yor process, which is widely used in the Bayesian nonparametric framework. By employing an infinite-dimensional model induced by the Pitman–Yor process, the proposed method accounts for the contribution of unobserved species and enables stable entropy estimation. Although the model distribution differs from the true distribution $\bm{p}$, simulation results show that our method performs well in estimating $H$, particularly when $K$ is large. Moreover, it performs comparably to the aforementioned conventional estimators when NN is large.
@@ 3.1 Dirichlet–Pitman–Yor mixture model
Given a random sample of size NN, let $T$ denote the number of observed species, and $n_{i}$ the number of observations of species $i$ for $i=1,\ldots,T$. We write $\mathbf{y}=(n_{1},\dots,n_{T})^{\top}$. Then, we have $N=\sum_{i=1}^{T}n_{i}$ by definition. Archer et al. [1] introduced the following mixture model for constructing a probability vector based on the Pitman–Yor process.
@@
Under the DPYM, given the observed frequency $\mathbf{y}=(n_{1},\dots,n_{T})^{\top}$ of NN individuals, the predictive distribution for a new observation $Z$ is given by
@@ 3.2 Data-driven parameter selection
Let $\bm{p}$ be a $K$-dimensional probability vector, and suppose that a random sample of size NN is drawn from $\bm{p}$, in which $T$ distinct species are observed. Let $\mathbf{y}=(n_{1},\ldots,n_{T})$ denote the corresponding observation counts. Then, the cross entropy $H(\bm{p},\bm{q})$ between $\bm{p}$ and $\bm{q}$ defined by (3.4) satisfies
@@ 3.2.3 Selection in the large-sample regime
The advantage of using DPYM is assigning positive probabilities to unobserved species. However, when the sample size NN is sufficiently large, we typically have $T=K$, meaning that all species have been observed. In this case, the terms corresponding to unobserved probabilities, $C_{0}$ and $C_{1}$, vanish. Furthermore, in this subsection, we assume that the indices of the observed counts $(n_{1},\ldots,n_{K})$ are aligned with the probability vector $\bm{p}=(p_{1},\ldots,p_{K})$. The upper bound function then reduces to
@@
which depends solely on NN and thus provides no meaningful information for determining $(d,\alpha)$. Hence, when NN is large, we adopt a strategy of directly minimizing the cross entropy to select the hyperparameters $(d,\alpha)$ rather than relying on the upper bound function. The cross entropy $H(\bm{p},\bm{q})$ can be rewritten as
@@ 4 Proposed estimator
Based on the preparation above, we summarize the construction of the proposed estimator $\hat{H}$. Let $T$ denote the number of observed species and $n_{i}$ the count of the $i$-th observed species for $i=1,\ldots,T$, with $\mathbf{y}=(n_{1},\dots,n_{T})^{\top}$. The sample size NN then satisfies $N=\sum_{i=1}^{T}n_{i}$.
@@
In addition, we consider the boundary case by minimizing (4.1) with respect to $\alpha$ when $d=0$ or when $d$ is sufficiently close to 11. When $d=0$, the corresponding value of $\alpha$ is
@@
When $d$ is sufficiently close to 11, i.e., $d=1-\epsilon$ for small $\epsilon>0$, the corresponding value is
@@ 5.1 Regularly varying function and regularly varying distribution
The proposed method is designed for settings in which the sample size NN is relatively small compared to the true number of species $K$. To ensure that the method retains a certain level of accuracy when NN is sufficiently large, we establish its consistency in the large-sample regime. The concepts of regularly varying functions and regularly varying distributions are used to prove this result, which are needed to verify the finiteness of the entropy. We summarize the relevant fundamentals in this subsection. For further details on regularly varying distributions, see Nair et al. [19], Resnick [23], and Bingham et al. [5].
@@ 5.2 Finiteness of the Shannon entropy of regularly varying distributions
where AA is a constant. Hence, finiteness of the first term on the right-hand side implies finiteness of $H(\bm{p})$. The divergence case follows from an analogous argument. This completes the proof. ∎
@@
where AA is a constant. Therefore, it holds that
@@ 6 Consistency of the proposed entropy estimator
By the law of large numbers, we have $\mathsf{P}(N^{-1}\mathbf{x}_{N}=\bm{p})\to 1$. Since $p_{k}>0$ for all $k\in\{1,\ldots,K\}$, it follows that $\mathsf{P}(m_{N,1}=0)\to 1$, and consequently $\mathsf{P}((\hat{d}_{N},\hat{\alpha}_{N})=(d_{0},\alpha_{0}))\to 1$. Moreover, $\mathsf{P}(T_{N}=K)\to 1$. Define the events
@@ 7.2 Simulation setup
Let us describe simulation settings. The dimension $K$ of the true probability vector $\bm{p}$ is fixed at $5000$, and the sample size NN is set to one of the following values: 10, 30, 50, 100, 200, 400, 600, 800, 1000, 3000, 5000, 8000, 10000, 20000. The probability vector $\bm{p}$ is generated according to the following five scenarios:
@@ 7.3 Summary of results from simulations
  • When NN is extremely small relative to the dimension $K$ of $\bm{p}$, the proposed estimator consistently outperforms the other estimators across all scenarios.
@@
  • Even when NN is sufficiently large, the proposed estimator achieves comparable error to the other methods, demonstrating stable estimation.
  • As the distribution becomes closer to uniform, estimation error of our proposed estimator tends to increase when NN is on the same order as the dimension $K$ of $\bm{p}$.

### v2

@@ 1 Introduction
In this study, we consider the problem of estimating $H$ in the setting where both $K$ and $\bm{p}$ are unknown, as in [1, 13]. Let $T$ denote the number of observed species in the random sample and let $n_{1},\ldots,n_{T}$ denote their observed frequencies. When the sample size $N$ is sufficiently large relative to $K$, it is reasonable to assume that $T=K$. In this case, the simplest and most widely used approach is to estimate $p_{1},\ldots,p_{T}$ by maximum likelihood and substitute the resulting estimates into (1.1). Specifically, the maximum likelihood estimator of each probability mass is given by
@@
In the literature, alternative estimators have been proposed, including the Miller–Madow estimator [18] and the Chao–Shen estimator [7]. However, it has been indicated that these methods tend to perform poorly when the sample size $N$ is not sufficiently large relative to the true number of species $K$ [20, 24]. To address this issue, Hausser and Strimmer [12] proposed a shrinkage estimator for the individual probability masses toward $1/K$ under the assumption that $K$ is known, and demonstrated that the resulting estimator exhibits favorable properties even for small sample sizes. In contrast, we consider the more challenging setting in which the sample size $N$ is smaller than the true number of species $K$, and $K$ is unknown. Under this regime, we propose a new estimator of $H$. The proposed approach is based on the Pitman–Yor process, which is widely used in the Bayesian nonparametric framework. By employing an infinite-dimensional model induced by the Pitman–Yor process, the proposed method accounts for the contribution of unobserved species and enables stable entropy estimation. Although the model distribution differs from the true distribution $\bm{p}$, simulation results show that our method performs well in estimating $H$, particularly when $K$ is large. Moreover, it performs comparably to the aforementioned conventional estimators when $N$ is large.
@@ 3.1 Dirichlet–Pitman–Yor mixture model
Given a random sample of size $N$, let $T$ denote the number of observed species, and $n_{i}$ the number of observations of species $i$ for $i=1,\ldots,T$. We write $\mathbf{y}=(n_{1},\dots,n_{T})^{\top}$. Then, we have $N=\sum_{i=1}^{T}n_{i}$ by definition. Archer et al. [1] introduced the following mixture model for constructing a probability vector based on the Pitman–Yor process.
@@
Under the DPYM, given the observed frequency $\mathbf{y}=(n_{1},\dots,n_{T})^{\top}$ of $N$ individuals, the predictive distribution for a new observation $Z$ is given by
@@ 3.2 Data-driven parameter selection
Let $\bm{p}$ be a $K$-dimensional probability vector, and suppose that a random sample of size $N$ is drawn from $\bm{p}$, in which $T$ distinct species are observed. Let $\mathbf{y}=(n_{1},\ldots,n_{T})$ denote the corresponding observation counts. Then, the cross entropy $H(\bm{p},\bm{q})$ between $\bm{p}$ and $\bm{q}$ defined by (3.4) satisfies
@@ 3.2.3 Selection in the large-sample regime
The advantage of using DPYM is assigning positive probabilities to unobserved species. However, when the sample size $N$ is sufficiently large, we typically have $T=K$, meaning that all species have been observed. In this case, the terms corresponding to unobserved probabilities, $C_{0}$ and $C_{1}$, vanish. Furthermore, in this subsection, we assume that the indices of the observed counts $(n_{1},\ldots,n_{K})$ are aligned with the probability vector $\bm{p}=(p_{1},\ldots,p_{K})$. The upper bound function then reduces to
@@
which depends solely on $N$ and thus provides no meaningful information for determining $(d,\alpha)$. Hence, when $N$ is large, we adopt a strategy of directly minimizing the cross entropy to select the hyperparameters $(d,\alpha)$ rather than relying on the upper bound function. The cross entropy $H(\bm{p},\bm{q})$ can be rewritten as
@@ 4 Proposed estimator
Based on the preparation above, we summarize the construction of the proposed estimator $\hat{H}$. Let $T$ denote the number of observed species and $n_{i}$ the count of the $i$-th observed species for $i=1,\ldots,T$, with $\mathbf{y}=(n_{1},\dots,n_{T})^{\top}$. The sample size $N$ then satisfies $N=\sum_{i=1}^{T}n_{i}$.
@@
In addition, we consider the boundary case by minimizing (4.1) with respect to $\alpha$ when $d=0$ or when $d$ is sufficiently close to $1$. When $d=0$, the corresponding value of $\alpha$ is
@@
When $d$ is sufficiently close to $1$, i.e., $d=1-\epsilon$ for small $\epsilon>0$, the corresponding value is
@@ 5.1 Regularly varying function and regularly varying distribution
The proposed method is designed for settings in which the sample size $N$ is relatively small compared to the true number of species $K$. To ensure that the method retains a certain level of accuracy when $N$ is sufficiently large, we establish its consistency in the large-sample regime. The concepts of regularly varying functions and regularly varying distributions are used to prove this result, which are needed to verify the finiteness of the entropy. We summarize the relevant fundamentals in this subsection. For further details on regularly varying distributions, see Nair et al. [19], Resnick [23], and Bingham et al. [5].
@@ 5.2 Finiteness of the Shannon entropy of regularly varying distributions
where $A$ is a constant. Hence, finiteness of the first term on the right-hand side implies finiteness of $H(\bm{p})$. The divergence case follows from an analogous argument. This completes the proof. ∎
@@
where $A$ is a constant. Therefore, it holds that
@@ 6 Consistency of the proposed entropy estimator
By the law of large numbers, we have $N^{-1}\mathbf{x}_{N}\to^{p}\bm{p}$. Since $p_{k}>0$ for all $k\in\{1,\ldots,K\}$, it follows that $\mathsf{P}(m_{N,1}=0)\to 1$, and consequently $\mathsf{P}((\hat{d}_{N},\hat{\alpha}_{N})=(d_{0},\alpha_{0}))\to 1$. Moreover, $\mathsf{P}(T_{N}=K)\to 1$. Define the events
@@ 7.2 Simulation setup
Let us describe simulation settings. The dimension $K$ of the true probability vector $\bm{p}$ is fixed at $5000$, and the sample size $N$ is set to one of the following values: 10, 30, 50, 100, 200, 400, 600, 800, 1000, 3000, 5000, 8000, 10000, 20000. The probability vector $\bm{p}$ is generated according to the following five scenarios:
@@ 7.3 Summary of results from simulations
  • When $N$ is extremely small relative to the dimension $K$ of $\bm{p}$, the proposed estimator consistently outperforms the other estimators across all scenarios.
@@
  • Even when $N$ is sufficiently large, the proposed estimator achieves comparable error to the other methods, demonstrating stable estimation.
  • As the distribution becomes closer to uniform, estimation error of our proposed estimator tends to increase when $N$ is on the same order as the dimension $K$ of $\bm{p}$.